Estimate the minimum potential difference needed to reduce Al2O3 at 5000C. The free energy change for the decomposition reaction is:


2/3 Al2O3----------------->4/3Al +O2
























?G=+960kJ

Grade:12

1 Answers

askiitian.expert- chandra sekhar
14 years ago

Hi Belani,

2/3 Al2O3----------------->4/3Al +O2


ΔG= +960 kJ for 2/3moles

But ΔG=-nFE

E = ΔG/(-nF)

F = 96500 C mol-1

E = 960×1000/(- 2/3 × 96500) = -14.9 V

This implies that the minimum potential difference needed is 14.9 V

 

askiitian.expert- chandra sekhar